# R&D Networks under Heterogeneous Firm Productivities

### v1

@@ 3 Simulation Exercises for Stability and Welfare
The analysis so far shows that the complete network need not remain stable when productivities are heterogeneous, unlike goyal2001r, where they also show that this structure is not welfare-maximizing in the sense that less connectivity is preferred in the social optimum. Building on these insights, in this section, we extend the analysis to investigate both the stability and welfare properties of alternative network structures in the heterogeneous-productivity setting.
Due to the intractable nature of the problem discussed earlier, we are unable to fully characterize the set of pairwise stable networks in general. Under homogeneity, goyal2001r encounter similar difficulties in determining whether the complete network is the unique stable (symmetric) configuration, and show this is the case by restricting the problem to $n=4$ firms. [^5] Readers are referred to pp. 696 of goyal2001r for a detailed discussion on the intractability of the network formation problem in this class of models. This challenge becomes much more pronounced under heterogeneity. Additionally, this intractability extends to making comparisons of equilibrium effort and profit levels across different network structures. Even after deriving closed-form solutions for efforts via Equation (7) in simple network configurations —found by inverting matrix $\mathbf{A}(\mathbf{G})$ that depends on the network— comparing levels of efforts across structures requires comparing rational functions of large degree in their denominators, which is again unfeasible even for small number of firms $n$.
@@
By using computer-assisted analysis, first, we fully characterize the set of stable networks for $n=4$ in this setting. [^6] For the full implementation of the computer-assisted setup, see the supplementary material. Crucially, the analysis reveals that differences in productivity levels and the distribution of high- and low-type firms within the economy can sustain pairwise stable networks different from the complete network.
Our first simulation documents a new fact: when introducing heterogeneity in firm R&D productivities, the positive assortative (PA) network $\mathbf{G}^{PA}$ —in which firms fully connect only to the others of their same type— becomes a pairwise stable configuration. In a two-types setting, a PA network is the one in which high- (low-) type firms are linked to every other high- (low-) type firms ($G_{ij}^{PA}=G_{ji}^{PA}=1$ if $i$ and $j$ are of the same type, otherwise $G_{ij}^{PA}=G_{ji}^{PA}=0$). This structure leads to two fully connected components of size $\rho\cdot n$ and $(1-\rho)\cdot n$ composed solely of firms of high- and low-type firms, respectively.
@@
Our results show that, for any fixed R&D effort cost coefficient $\phi$, there exist threshold levels $\underline{\theta}$ (shown as the green dashed line) and $\overline{\theta}$ (shown as the blue dashed line) with $\underline{\theta}<\overline{\theta}$. The complete network is stable for $\underline{\theta}\leq\theta<1$, and the PA network is stable for $0<\theta\leq\overline{\theta}$. These together imply both networks are stable in the interim region $\underline{\theta}\leq\theta\leq\overline{\theta}$. In addition, both thresholds are decreasing in $\phi$, meaning that the stability area spanned by the complete/PA network is increasing/decreasing in R&D cost.
Furthermore, while the PA and complete networks span the entire parameter space for $\theta$, there is one other configuration that emerges as stable for certain parameter values, which is shown in the middle right panel of Figure 3. This network is an intermediate between PA and complete, in the sense that it is denser/sparser than the PA/complete network through adding/severing links from a high-type firm to all low-type firms. Lastly, this full characterization shows that for sufficiently low/high $\theta$, the PA/complete network is the unique stable structure.
@@ 3.1 Welfare Comparison
Following goyal2001r, we define welfare as the total surplus generated in the economy by summing up producer and consumer surplus. [^7] For an alternative approach based on a utility function over consumption, see hsieh2024endogenous. Producer surplus ($PS$) is the sum of profits earned by all firms in the network, given by $PS\coloneqq\sum_{i=1}^{n}{\pi}^{*}_{i},$ and consumer surplus ($CS$) is given by $CS\coloneqq(1/2)\left(\sum_{i=1}^{n}{q}^{*}_{i}\right)^{2},$ as derived from the linear demand specification $p(\mathbf{q})$. Adding up these components, total welfare of the given R&D network structure $\mathbf{G}$ is given by:
@@ 3.2 Crowding-out Effect
Figure 5 provides an additional insight: as $\rho$ rises, while the stability region of the complete network remains unchanged, this is not the case for the PA structure. Instead, the PA stability region shrinks as $\rho$ increases. Formally, the upper threshold $\overline{\theta}$, which ensures the stability of the PA network for any $\theta\leq\overline{\theta}$, decreases in the fraction of high-productivity firms, while the lower threshold $\underline{\theta}$, guaranteeing the stability of the complete network for any $\theta\geq\underline{\theta}$, remains unaffected. [^9] For a comparison of how the stability regions of the PA and complete structures evolve under different settings, see Figure 9 in Appendix B.
As shown earlier in this Section (and illustrated in Figures 3 and 4), the PA and complete network configurations are the main pairwise-stable structures under heterogeneity: together, they span almost the entire parameter space, meaning that when one structure is not stable, the other is typically stable. Therefore, understanding the crowding-out patterns in the PA structure becomes especially relevant in alternative economies where the PA structure is stable.
@@
This implies that as $\rho$ increases from 0 to 1, the number of links decreases until reaching a minimum at $\rho=0.5$, symmetrically increasing thereafter. As goyal2001r show, welfare in symmetric networks under the homogeneous setting exhibits an inverted-U relationship with the number of links. This naturally suggests that the crowding-out effect we observe in the PA structure under heterogeneity may stem from a generalized version of their result, where welfare displays an inverted-U shape in the network’s overall connectivity.
@@
For symmetric networks with homogeneous firms, goyal2001r show that welfare is concave in the average degree of the network, implying that welfare is maximized at an intermediate level of connectivity—a network configuration between the empty and complete. In our heterogeneous setting, we find a similar and more general connectivity effect, based on simulations with randomly generated networks instead of the more restricted symmetric network case. In Figure 6, we depict the inverted U-shaped relationship between the network density and welfare for $n=10$ firms, where, for a given total number of links (on the $x$-axis), we average welfare over 1000 replications of networks where the links are randomly assigned.
@@
Together, Figures 6 and 7 indicate that the performance of the PA structure is highly non-monotonic relative to the average benchmark. At extreme values of $\rho$ (either low or high), the welfare in PA network is lower relative to the average welfare in random networks with equal link density. Conversely, at intermediate levels of $\rho$, the PA structure generates a significant welfare premium. This premium, however, diminishes as the productivity gap closes (i.e., as $\theta$ increases). The network structure advantage at intermediate $\rho$ is so pronounced in large productivity gap settings that it can override the effect of lower intrinsic productivity. For instance, at $\rho=0.5$, the PA network with a large productivity gap ($\theta=0.1$) achieves a welfare level exceeding even the maximum average welfare of random networks with a significantly smaller productivity gap ($\theta=0.5$).
@@ 4 Conclusion
In this paper, we study R&D network formation with heterogeneous productivities, generalizing the homogeneous framework in goyal2001r. Firms differ in their cost of doing R&D, modeled as a reduction in the marginal cost of production in a Cournot competition setup.
Under such heterogeneity, first, we provide some comparative statics results on R&D efforts and profits of firms. The findings show connectivity-based relative efforts and profits: when a low and a high-productivity firm (having a symmetric position in the R&D network) gets connected, the relative gain of the less productive firm is higher than the more productive firm. Moreover, in terms of efforts, when two firms (with symmetric network positions) are compared, the one with higher productivity has more R&D effort, and hence total R&D cost, than the one with lower productivity.
These benchmark comparisons show that low-productivity firms highly benefit from being connected to high-productivity firms, but the reverse does not necessarily hold, having explicit implications for link formation incentives. In particular, the complete network— that is pairwise stable in a homogeneous setting—ceases to be stable when the productivity gap is sufficiently large.
Our simulation results illustrates a stability feature: the positive assortative (PA) network structure, which is a productivity-level-based clustered network, and the complete network structure together almost covers the entire stability space, meaning that when one of these two structures is not stable, the other typically is. Therefore, understanding the welfare and stability implications of the PA structure is crucial. This alternative configuration has unique implications for welfare, defined as the aggregate consumer and producer surplus in the economy.
Our simulations show that the PA structure yields higher welfare in comparison to the complete structure. Furthermore, we uncover a counterintuitive effect: a higher average productivity in the economy could potentially lead to a lower welfare due to (i) emergence of alternative stable structures and (ii) crowding-out effect of high-productivity firms. Finally, we compare the welfare in the PA and complete networks with the average welfare in random networks in our simulations, isolating the welfare effects of the network structure from the welfare effects of network density measured by the number of links, providing further insights into how the connectivity structure together with productivity distribution and network density shape the welfare under heterogeneity.
Taken together, these findings suggest that economies with higher average firm-level R&D productivity—either through higher productivity of low-type firms or through a larger fraction of high-type firms—may exhibit lower welfare compared to economies with lower average R&D productivity. Therefore, policies aimed at enhancing firm level R&D productivities should account for their impact on endogenous R&D alliances and effort, as endogenous responses to productivity changes may offset or even reverse the intended welfare gains.
@@ Appendix A Proofs
We use the Gershgorin’s theorem: a matrix is nonsingular if every column is strictly diagonally dominant, i.e.
$|A_{ii}|\;>\;\sum_{j\neq i}|A_{ji}|\qquad\text{for all }i.$
For $i$ we have
$\sum_{j\neq i}|A_{ji}|$ $=\sum_{j\neq i}\left|\theta_{j}\big\{(n+1)\,G_{ij}-[1+d_{j}(\mathbf{G})]\big\}\right|$
$=\sum_{j\neq i}\theta_{j}\big|(n+1)\,G_{ij}-d_{j}(\mathbf{G})-1\big|$
$=\sum_{j\in\mathcal{N}_{i}(\mathbf{G})}\theta_{j}\big|n-d_{j}(\mathbf{G})\big|\;+\;\sum_{j\in\mathcal{N}_{-i}(\mathbf{G})}\theta_{j}\big|-d_{j}(\mathbf{G})-1\big|,$
where $\mathcal{N}_{i}(\mathbf{G})$ denotes the neighbors of $i$ and $\mathcal{N}_{-i}(\mathbf{G})$ its non-neighbors.
Bounding each term by its maximum gives
$\sum_{j\neq i}|A_{ji}|<d_{i}(\mathbf{G})\max_{j\in\mathcal{N}_{i}(\mathbf{G})}\!\big[n-d_{j}(\mathbf{G})\big]+(n-1-d_{i}(\mathbf{G}))\max_{j\in\mathcal{N}_{-i}(\mathbf{G})}\!\big[d_{j}(\mathbf{G})+1\big].$
If $j\in\mathcal{N}_{i}(\mathbf{G})$, then $d_{j}(\mathbf{G})\geq 1$ so
$\max_{j\in\mathcal{N}_{i}(\mathbf{G})}\!\big[n-d_{j}(\mathbf{G})\big]\;\leq\;n-1.$
If $j\in\mathcal{N}_{-i}(\mathbf{G})$, then $d_{j}(\mathbf{G})\leq n-2$ so
$\max_{j\in\mathcal{N}_{-i}(\mathbf{G})}\!\big(d_{j}(\mathbf{G})+1\big)\;\leq\;n-1.$
Therefore,
$\sum_{j\neq i}|A_{ji}|<d_{i}(\mathbf{G})\,(n-1)+[n-1-d_{i}(\mathbf{G})]\,\big(n-1\big)=(n-1)^{2}.$ (1)
Column $i$ is strictly diagonally dominant if
$\frac{(n+1)^{2}\phi}{\theta_{i}[n-d_{i}(\mathbf{G})]}-\theta_{i}[n-d_{i}(\mathbf{G})]>(n-1)^{2}.$
Rearranging yields the sufficient bound
$\phi>\theta_{i}\,[n-d_{i}(\mathbf{G})]\left(\frac{n-1}{n+2}\right)^{2}+\theta_{i}^{2}\,\left(\frac{n-d_{i}(\mathbf{G})}{n+1}\right)^{2}.$ (2)
To obtain a uniform sufficient condition, set $\theta_{i}\leq\theta_{\max}=1$ and $n-d_{i}(\mathbf{G})\leq n$. Therefore
@@
Thus if $\phi$ satisfies the above inequality, every column of $A$ is strictly diagonally dominant, and by Gershgorin’s theorem the matrix $A$ is invertible.
@@
We can rewrite the FOC in equation (6) in equilibrium as
$\left[\frac{\phi}{\theta_{i}\,\eta_{i}(\mathbf{G})}-\theta_{i}\,\eta_{i}(\mathbf{G})\right]{e}^{*}_{i}(\mathbf{G})=\frac{\alpha-\bar{c}}{n+1}+\sum_{l\in\mathcal{N}_{i}(\mathbf{G})}\theta_{l}\,\eta_{l}(\mathbf{G})\,{e}^{*}_{l}(\mathbf{G})-\sum_{k\in\mathcal{N}_{-i}(\mathbf{G})}[1-\eta_{k}(\mathbf{G})]\,\theta_{k}\,{e}^{*}_{k}(\mathbf{G}).$
Lets define $\mathcal{N}_{i}^{(-j)}(\mathbf{G})\coloneqq\{k\;|\;G_{ik}=1,k\neq i,j\}$, and $\mathcal{N}_{-i}^{(-j)}(\mathbf{G})\coloneqq\{k\;|\;G_{ik}=0,k\neq i,j\}$ to be the set on neighbors and non-neighbors of firm $i$ in network $\mathbf{G}$ excluding the firm $j$. Thus for every pair of firms $i$ and $j$ we can write:
$\left[\frac{\phi}{\theta_{i}\,\eta_{i}(\mathbf{G})}-\theta_{i}\,\eta_{i}(\mathbf{G})\right]{e}^{*}_{i}(\mathbf{G})-\left[\eta_{j}(\mathbf{G})-1+G_{ij}\right]\,\theta_{j}{e}^{*}_{j}(\mathbf{G})=\frac{\alpha-\bar{c}}{n+1}\\ +\sum_{l\in\mathcal{N}^{(-j)}_{i}(\mathbf{G})}\theta_{l}\,\eta_{l}(\mathbf{G})\,{e}^{*}_{l}(\mathbf{G})-\sum_{k\in\mathcal{N}^{(-j)}_{-i}(\mathbf{G})}[1-\eta_{k}(\mathbf{G})]\,\theta_{k}\,{e}^{*}_{k}(\mathbf{G}).$
From the symmetric position definition in Definition 1 we know that if two firm $i$ and $j$ have a symmetric position with respect to a network means:
$\mathcal{N}_{i}^{(-j)}(\mathbf{G})=\mathcal{N}_{j}^{(-i)}(\mathbf{G}),\quad\text{and}\quad\mathcal{N}_{-i}^{(-j)}(\mathbf{G})=\mathcal{N}_{-j}^{(-i)}(\mathbf{G}).$
Therefore the right hand side of the rewritten version of equilibrium FOCs for the symmetric firms $i$ and $j$ are equal. Therefore:
$\left[\frac{\phi}{\theta_{i}\,\eta_{i}(\mathbf{G})}-\theta_{i}\,\eta_{i}(\mathbf{G})\right]{e}^{*}_{i}(\mathbf{G})-\left[\eta_{j}(\mathbf{G})-1+G_{ij}\right]\,\theta_{j}{e}^{*}_{j}(\mathbf{G})\\ =\\ \left[\frac{\phi}{\theta_{j}\,\eta_{j}(\mathbf{G})}-\theta_{j}\,\eta_{j}(\mathbf{G})\right]{e}^{*}_{j}(\mathbf{G})-\left[\eta_{i}(\mathbf{G})-1+G_{ji}\right]\,\theta_{i}{e}^{*}_{i}(\mathbf{G}).$ (14)
Rearranging the equation, we get:
@@
Finally, substituting this into the equilibrium profit equation (9) we get:
@@
Showing that ${\pi}^{*}_{i}(\mathbf{G}_{+ij})<{\pi}^{*}_{j}(\mathbf{G}_{+ij})$ is straightforward as we have:
@@
Since $\eta^{2}_{i}(\mathbf{G}_{+ij})=\eta^{2}_{j}(\mathbf{G}_{+ij})$ due to the symmetric position assumption, the term $\phi-\theta_{i}^{2}\;\eta^{2}_{i}(\mathbf{G}_{+ij})>\phi-\theta_{j}^{2}\;\eta^{2}_{j}(\mathbf{G}_{+ij})$, and ${\pi}^{*}_{i}(\mathbf{G}_{+ij})/{\pi}^{*}_{j}(\mathbf{G}_{+ij})<1$. Now for when $G_{ij}=0$ we have:
@@
where we define $f(x):=\left[\phi-x\,\eta_{i}^{2}(\mathbf{G}_{-ij})\right]/\left[\phi-x\,\eta_{i}(\mathbf{G}_{-ij})\right]$. The last fraction is greater that 1, as $\theta_{i}>\theta_{j}$. Now we only need to show that $f(x)$ is an strictly increasing function, which makes the first two fraction greater that 1 as well and results in ${\pi}^{*}_{i}(\mathbf{G}_{-ij})>{\pi}^{*}_{i}(\mathbf{G}_{-ij})$.
Taking the derivative of $f$ with respect to $x$, and by using the term $\zeta\coloneqq\eta_{i}(\mathbf{G}_{-ij})$ for convinience, we have:
$\frac{\partial f(x)}{\partial x}$ $=\frac{\partial}{\partial x}\left(\frac{\phi-\zeta^{2}\,x}{\phi-\zeta\,x}\right)$
$=\frac{-\zeta^{2}\,(\phi-\zeta\,x)+\zeta\,(\phi-\zeta^{2}\,x)}{(\phi-\zeta\,x)^{2}}$
$=\frac{\zeta\,(\phi-\zeta\,\phi)}{(\phi-\zeta\,x)^{2}}=\frac{\zeta\,(1-\zeta)\,\phi}{(\phi-\zeta\,x)^{2}}.$
Since $\zeta=\eta_{i}(\mathbf{G}_{-ij})=\eta_{j}(\mathbf{G}_{-ij})\in\left[\frac{2}{n+1},\frac{n}{n+1}\right]$, we have $f^{\prime}(x)>0$, meaning that $f(x)$ is an strictly increasing function, and therefore:
${\pi}^{*}_{i}(\mathbf{G}_{-ij})>{\pi}^{*}_{i}(\mathbf{G}_{-ij}).$

### v2

@@ 3 Simulation exercises for stability and welfare
The analysis so far shows that the complete network need not remain stable when productivities are heterogeneous, unlike Goyal and Moraga-González (2001), where they also show that this structure is not welfare-maximizing in the sense that less connectivity is preferred in the social optimum. Building on these insights, in this section, we extend the analysis to investigate both the stability and welfare properties of alternative network structures in the heterogeneous-productivity setting.
Due to the intractable nature of the problem discussed earlier, we are unable to fully characterize the set of pairwise stable networks in general. Under homogeneity, Goyal and Moraga-González (2001) encounter similar difficulties in determining whether the complete network is the unique stable (symmetric) configuration, and show this is the case by restricting the problem to $n=4$ firms. [^5] Readers are referred to pp. 696 of Goyal and Moraga-González (2001) for a detailed discussion on the intractability of the network formation problem in this class of models. This challenge becomes much more pronounced under heterogeneity. Additionally, this intractability extends to making comparisons of equilibrium effort and profit levels across different network structures. Even for networks in which the matrix $\mathbf{A}(\mathbf{G})$ can be inverted to provide a closed-form solutions for efforts via Equation (7), comparing effort levels across structures requires comparing rational functions of large degree, which is again unfeasible even for small number of firms $n$.
@@
We begin by fully characterizing the set of stable networks for $n=4$ firms, by using computer-assisted derivations for the closed form solutions. [^6] The full implementation of the computer-assisted derivations are in the Supplementary Appendix. Crucially, the analysis reveals that differences in productivity levels and the distribution of high- and low-type firms within the economy can sustain pairwise stable networks different from the complete network.
Our first simulation documents a new fact: when introducing heterogeneity in firm R&D productivities, the positive assortative (PA) network $\mathbf{G}^{PA}$ —in which firms fully connect only to the others of their same type— becomes a pairwise stable configuration. In a two-types setting, a PA network is the one in which high- (low-) type firms are only linked to other high- (low-) type firms ($G_{ij}^{PA}=G_{ji}^{PA}=1$ if $i$ and $j$ are of the same type, otherwise $G_{ij}^{PA}=G_{ji}^{PA}=0$). This structure leads to two fully connected components of size $\rho\cdot n$ and $(1-\rho)\cdot n$, composed solely of firms of high- and low-type firms, respectively.
@@
Our results show that, for any fixed R&D effort cost coefficient $\phi$, there exist threshold levels $\underline{\theta}$ (shown as the green dashed line) and $\overline{\theta}$ (shown as the blue dashed line) with $\underline{\theta}<\overline{\theta}$. The complete network is stable for $\underline{\theta}\leq\theta<1$, and the PA network is stable for $0<\theta\leq\overline{\theta}$. Together, these facts imply both networks are stable in the interim region $\underline{\theta}\leq\theta\leq\overline{\theta}$. In addition, both thresholds are decreasing in $\phi$, meaning that the stability area spanned by the complete/PA network is increasing/decreasing in R&D cost.
Furthermore, while the PA and complete networks span the entire parameter space for $\theta$, there is one other configuration that emerges as stable for certain parameter values, which is shown in the middle right panel of Figure 2. This network is an intermediate between PA and complete, in the sense that it is denser/sparser than the PA/complete network through adding/severing links from a high-type firm to all low-type firms. Lastly, this characterization shows that for sufficiently low/high $\theta$, the PA/complete network is the unique stable structure.
@@ 3.1 Welfare Comparison
Following Goyal and Moraga-González (2001), we define welfare as the total surplus generated in the economy by adding producer and consumer surplus. [^7] For an alternative approach based on a utility function over consumption, see Hsieh et al. (2025). Producer surplus ($PS$) is the sum of profits earned by all firms in the network, given by $PS\coloneqq\sum_{i=1}^{n}{\pi}^{*}_{i},$ and consumer surplus ($CS$) is given by $CS\coloneqq(1/2)\left(\sum_{i=1}^{n}{q}^{*}_{i}\right)^{2},$ as derived from the linear demand specification $p(\mathbf{q})$. Adding up these components, the total welfare of any given R&D network structure $\mathbf{G}$ is given by:
@@ 3.2 Crowding-out Effect
Figure 4 provides an additional insight: as $\rho$ rises, while the stability region of the complete network remains unchanged, this is not the case for the PA structure. Instead, the PA stability region shrinks as $\rho$ increases. Formally, the upper threshold $\overline{\theta}$, which ensures the stability of the PA network for any $\theta\leq\overline{\theta}$, decreases in the fraction of high-productivity firms, while $\underline{\theta}$, guaranteeing the stability of the complete network for any $\theta\geq\underline{\theta}$, remains unaffected. [^9] For a comparison of how the stability regions of the PA and complete structures evolve under different settings, see Figure 8 in Appendix B.
As shown earlier in this Section (and illustrated in Figures 2 and 3), the PA and complete network configurations are the main pairwise-stable structures under heterogeneity: together, they span almost the entire parameter space, meaning that when one structure is not stable, the other is typically stable. Therefore, understanding the crowding-out patterns in the PA structure becomes especially relevant in alternative economies where the PA structure is stable.
@@
This implies that as $\rho$ increases from 0 to 1, the number of links decreases until reaching a minimum at $\rho=0.5$, symmetrically increasing thereafter. As Goyal and Moraga-González (2001) show, welfare in symmetric networks under the homogeneous setting exhibits an inverted-U relationship with the number of links. This naturally suggests that the crowding-out effect we observe in the PA structure under heterogeneity may stem from a generalized version of their result, where welfare displays an inverted-U shape in the network’s overall connectivity.
@@
For symmetric networks with homogeneous firms, Goyal and Moraga-González (2001) show that welfare is concave in the average degree of the network, implying that welfare is maximized at an intermediate level of connectivity—a network configuration between the empty and complete. In our heterogeneous setting, we find a similar and more general connectivity effect, based on simulations with randomly generated networks instead of the more restricted symmetric network case. In Figure 5, we depict the inverted U-shaped relationship between the network density and welfare for $n=10$ firms, where, for a given total number of links (on the $x$-axis), we average welfare over 1000 replications of networks where the links are randomly assigned.
@@
Together, Figures 5 and 6 indicate that the performance of the PA structure is highly non-monotonic relative to the average benchmark. At extreme values of $\rho$ (either low or high), the welfare in PA network is lower relative to the average welfare in random networks with equal link density. Conversely, at intermediate levels of $\rho$, the PA structure generates a significant welfare premium. This premium, however, diminishes as the productivity gap closes (i.e., as $\theta$ increases). The network structure advantage at intermediate $\rho$ is so pronounced in large productivity gap settings that it can override the effect of lower intrinsic productivity. For instance, at $\rho=0.5$, the PA network with a large productivity gap ($\theta=0.1$) achieves a welfare level exceeding even the maximum average welfare of random networks with a significantly smaller productivity gap ($\theta=0.5$).
@@ 4 Conclusion
In this paper, we study R&D network formation with heterogeneous productivities, generalizing the homogeneous framework in Goyal and Moraga-González (2001). Firms differ in the potential benefit of their R&D activities and alliances, creating heterogeneous gains from own R&D efforts as well as from R&D externalities in a network setting.
Under such heterogeneity, first, we show that when a low- and a high-productivity firm (having a symmetric position in the R&D network) are connected, the relative gain of the less productive firm is higher than the more productive firm. Moreover, when such two firms are compared, the one with higher productivity exerts more R&D effort, and hence pays a higher R&D cost, than the one with lower productivity. These benchmark comparisons also imply that while low-productivity firms benefit from being connected to high-productivity firms, the reverse does not necessarily hold, creating heterogeneous incentives for link formation. We show the complete network—pairwise stable in a homogeneous setting—ceases to be stable when the productivity gap is sufficiently large.
By using computer-assisted closed form solutions, we next describe the full set of stable network structures in a 4-firms setting. Our findings illustrate a stability feature: the positive assortative (PA) network structure, which is a productivity-level-based clustered network, and the complete network structure together almost cover the entire stability space. This means that when one of these two structures is not stable, the other typically is. Therefore, understanding the welfare and stability implications of the PA structure is crucial. This alternative configuration has unique implications for welfare measured as the aggregate consumer and producer surplus in the economy.
The PA structure yields higher welfare in comparison to the complete structure, uncovering an R&D productivity puzzle: an economy with a higher average R&D productivity may exhibit lower welfare due to (i) emergence of alternative stable structures and (ii) crowding-out of high-productivity firms. Finally, we compare the welfare in the PA and complete networks with the average welfare in random networks using simulations, isolating the effects of network structure on welfare from the network density effect. This provides further insights into how the productivity distribution, connectivity structure and network density determine aggregate welfare under heterogeneity.
Altogether, our results highlight that R&D productivity fundamentally shapes the organization of innovation, with distinct implications for firm-level outcomes and aggregate welfare. These findings suggest that policies aimed at enhancing firm-level R&D productivities must account for their impact on endogenous R&D alliances and effort, as these responses may reverse the intended welfare gains.
@@ Appendix A Proofs
We first provide the proofs of the propositions presented in the paper. Statements and proofs of Lemmas used in the main results are provided at the end of the section.
@@
We use the Levy-Desplanques theorem, which is a consequence of Gershgorin’s disc theorem (Gershgorin, 1931): an $n\times n$ matrix $A$ is nonsingular if every column is strictly diagonally dominant, i.e. $|A_{ii}|>\sum_{j\neq i}|A_{ij}|$ for all $i=1,\ldots,n$. We will show that a bound on $\phi$ guarantees this property for $\mathbf{A}(\mathbf{G})$ as defined in (7). For any firm $i\in\mathcal{N}$, we have
$\sum_{j\neq i}|A_{ij}(\mathbf{G})|$ $=\sum_{j\neq i}\left|\theta_{j}\big\{(n+1)G_{ij}-[1+d_{j}(\mathbf{G})]\big\}\right|,$
$=\sum_{j\in\mathcal{N}_{i}(\mathbf{G})}\theta_{j}\big|n-d_{j}(\mathbf{G})\big|+\sum_{k\in\mathcal{N}_{-i}(\mathbf{G})}\theta_{k}\big|1+d_{k}(\mathbf{G})\big|.$
The cardinalities of the set of neighbours and non-neighbours are given by $|\mathcal{N}_{i}|=d_{i}(\mathbf{G})$ and $|\mathcal{N}_{-i}(\mathbf{G})|=n-1-d_{i}(\mathbf{G})$, respectively. For any neighbour $j\in\mathcal{N}_{i}(\mathbf{G})$, $1\leq d_{j}(\mathbf{G})\leq n-1$ so that $1\leq n-d_{j}(\mathbf{G})\leq n-1$, and similarly for non-neighbours $k\in\mathcal{N}_{-i}(\mathbf{G})$, $0\leq d_{k}(\mathbf{G})\leq n-2$ so that $1\leq 1+d_{k}(\mathbf{G})\leq n-1$. Additionally, recall $\theta_{i}\leq 1$ for all $i\in\mathcal{N}$. Bounding quantities using their maximum for all firms in each of the sets gives
$\sum_{j\neq i}|A_{ij}(\mathbf{G})|\leq d_{i}(\mathbf{G})\cdot(n-1)+[n-1-d_{i}(\mathbf{G})](n-1)=(n-1)^{2},\quad\text{for each }i\in\mathcal{N}.$
A sufficient condition that guarantees $\mathbf{A}(\mathbf{G})$ is strictly diagonally dominant is therefore
$\min_{i\in\mathcal{N}}|A_{ii}(\mathbf{G})|=\min_{i\in\mathcal{N}}\left|\frac{(n+1)^{2}\phi}{\theta_{i}[n-d_{i}(\mathbf{G})]}-\theta_{i}[n-d_{i}(\mathbf{G})]\right|>(n-1)^{2}.$
Using $\theta_{i}\leq 1$ and $n-d_{i}(\mathbf{G})\leq n$, we see that any $\phi\geq 1$ ensures $A_{ii}(\mathbf{G})>0$. Under these bounds, $\min_{i\in\mathcal{N}}|A_{ii}(\mathbf{G})|\geq[(n+1)^{2}\phi-n^{2}]/n$, and solving for $\phi$ yields the sufficient inequality
@@
The FOC in equation (6) for any firm $i\in\mathcal{N}$ can be written as
$\left[\frac{\phi}{\theta_{i}\eta_{i}(\mathbf{G})}-\theta_{i}\eta_{i}(\mathbf{G})\right]{e}^{*}_{i}(\mathbf{G})-\left[\eta_{j}(\mathbf{G})-1+G_{ij}\right]\theta_{j}{e}^{*}_{j}(\mathbf{G})=\frac{\alpha-\bar{c}}{n+1}\\ +\sum_{l\in\mathcal{N}_{i}(\mathbf{G})\setminus\{j\}}\theta_{l}\eta_{l}(\mathbf{G}){e}^{*}_{l}(\mathbf{G})-\sum_{k\in\mathcal{N}_{-i}(\mathbf{G})\setminus\{j\}}[1-\eta_{k}(\mathbf{G})]\theta_{k}{e}^{*}_{k}(\mathbf{G}).$
For any two firms $i$ and $j$ having a symmetric position (Definition 1), we have $\eta_{i}(\mathbf{G})=\eta_{j}(\mathbf{G})$ as well as
$\mathcal{N}_{i}(\mathbf{G})\setminus\{j\}=\mathcal{N}_{j}(\mathbf{G})\setminus\{i\}\quad\text{and}\quad\mathcal{N}_{-i}(\mathbf{G})\setminus\{j\}=\mathcal{N}_{-j}(\mathbf{G})\setminus\{i\}.$
Therefore, the right-hand side of the FOCs for firms $i$ and $j$ with symmetric position are equal, from which we can obtain
$\left[\frac{\phi}{\theta_{i}\eta_{i}(\mathbf{G})}-\theta_{i}\eta_{i}(\mathbf{G})\right]{e}^{*}_{i}(\mathbf{G})-\left[\eta_{j}(\mathbf{G})-1+G_{ij}\right]\theta_{j}{e}^{*}_{j}(\mathbf{G})=\\ \left[\frac{\phi}{\theta_{j}\eta_{j}(\mathbf{G})}-\theta_{j}\eta_{j}(\mathbf{G})\right]{e}^{*}_{j}(\mathbf{G})-\left[\eta_{i}(\mathbf{G})-1+G_{ji}\right]\theta_{i}{e}^{*}_{i}(\mathbf{G}).$
Rearranging the equation, we obtain the effort ratio between firms $i$ and $j$:
@@
Finally, substituting this into the equilibrium profit equation (9), we obtain the profit ratio:
@@
Showing that ${\pi}^{*}_{i}(\mathbf{G}_{+ij})<{\pi}^{*}_{j}(\mathbf{G}_{+ij})$ is straightforward, as we have:
@@
Due to the symmetric position assumption, $\eta^{2}_{i}(\mathbf{G}_{+ij})=\eta^{2}_{j}(\mathbf{G}_{+ij})$, so the term $\phi-\theta_{i}^{2}\eta^{2}_{i}(\mathbf{G}_{+ij})>\phi-\theta_{j}^{2}\eta^{2}_{j}(\mathbf{G}_{+ij})$ when $\theta_{i}>\theta_{j}$, implying ${\pi}^{*}_{i}(\mathbf{G}_{+ij})<{\pi}^{*}_{j}(\mathbf{G}_{+ij})$. Now, in the case when $G_{ij}=0$, we have
@@
where we define $f(x):=[\phi-x\cdot\eta_{i}^{2}(\mathbf{G}_{-ij})]/\left[\phi-x\cdot\eta_{i}(\mathbf{G}_{-ij})\right]$. The last fraction is greater than 1, as $\theta_{i}>\theta_{j}$. We now show that $f(x)$ is an strictly increasing function, which makes the product of the first two terms $f(\theta_{i}^{2})/f(\theta_{j}^{2})>1$, guaranteeing ${\pi}^{*}_{i}(\mathbf{G}_{-ij})>{\pi}^{*}_{i}(\mathbf{G}_{-ij})$. Taking the derivative of $f$ with respect to $x$, we have:
$\frac{df(x)}{dx}$ $=\frac{\phi\cdot\eta_{i}(\mathbf{G}_{-ij})[1-\eta_{i}(\mathbf{G}_{-ij})]}{[\phi-x\cdot\eta_{i}(\mathbf{G}_{-ij})]^{2}}.$
Since $\phi\geq 1$, $x\in(0,1]$, and $\eta_{i}(\mathbf{G}_{-ij})=\eta_{j}(\mathbf{G}_{-ij})\in[2/(n+1),n/(n+1)]\subset(0,1)$, we have $f^{\prime}(x)>0$. ∎